OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (–2, 1) and (4, 13)?

OpenStudy (anonymous):

(y-1) = 2(x+2)

OpenStudy (anonymous):

\[y _{2}-y _{1} = m(x _{2}-x _{1})\]

OpenStudy (anonymous):

that is the point-slope form equation :))